(x-5)over (x+7) equals (x+4)over (x+6)
Answer by mananth(16949)

(x-5)/(x+7) = (x+4)/ (x+6)
(x-5)(x+6)=(x+4)(x+7)
x^2+x-30=x^2 + 11x+28
x-11x=28-30
-10x=-2
x= -2/-10
x=1/5